Mir Status Report 27 June 1997

Date: Friday, June 27, 1997 // Mission Day 138/43

OPERATIONS / Public Affairs Support -- The two NASA PAOs at MCC-M continu e to respond to media inquiries regarding the health of Mike Foale, the i mpact of the current situation on NASA 5 science OPS, and the status of Mir. A Mir status report was forwarded to Houston and HQ PAO for review/dissemination. The PAOs are also endeavoring to protect the OPS team from the ever-present media at MCC-M.

Experiment Operations

TEF -- After Mir lost attitude control which required Soyuz maneuvering for correction, the TEF was requested to be powered off. It had only been powered up for approximately 8 hours. Temperature at the time of power on yesterday was +16 deg. C. The samples (2nd collection of the Sleep Experiment) will be stowed for return; however, there will be no future requests or priority placed on reactivating the TEF to preserve these samples.

Greenhouse -- Mike reported that the seed pods are looking pretty good a nd expressed his concern about losing them. Mike stated that we needed to think very hard about how long they had been in the dark and to think what we needed to do to preserve the results of the experiment. Mike was advised of the current plan for Greenhouse. The cabling required to power Greenhouse is being researched. If power can be supplied to the SVET, the Greenhouse lights will be turned on for a little while. Water will also be supplied to the plants (Program 3). The radiogram for this has been supplied by the Russian/US PI team and is in work. Mike then reported that the leaves had curled some and that there were no more flowers present. There were large seed pods and that they looked pretty normal in appearance and size per the diagrams. Mike stated that he was very interested in opening one up for a look.

MOST was advised by the Russian ground team that a manual ventilation of beetle would occur tomorrow and that attempted power up would occur on Monday, June 30. MOST is trying to verify, but believes that the plan is for a permanent power up.

Experiment Hardware : PHOTO/VIDEO - Mike reported that he has the 35 mm camera bag with accessories except for the 35 mm lens. The 35 mm lens, F -2, flash, and calibration for Greenhouse were all in Spektr. The 70 mm camera bags were left in Spektr; however, Mike had the two cameras (Hass elblad) with two 250 mm lens and one 100 mm lens, the IR and CPOS all sep arate in a box in Priroda. Mike also reported that he had the film chang eout bag.

Hardware status per Mike's calldown:

"in Spektr:
all of the SAMS Disks
all of the Sleep Experiment except Night Headband kits for the cosmonauts
Mir Supplemental Medical Kits
TEHOF Lids
Mike's PHK and PPK
Space Shuttle Sleeping Bag
Data Cassettes for FBI, Photos, and Film
Microbial Air Sampler
Solid Sorbent Air Sampler
Microbiological Monitoring Experiment
Treadmill Harness
Running Shoes (size 44 European, 11 1/2 American)
Sandals
GASMAP; TEHOF; Pathfinder
BDL ; UBC; COSS
Leaf Bags for Greenhouse
2 backup FBI hard-drives
backup drives for MIPS and spare PCMCIA cards
Cognilab disks
F-2 Camera used for Greenhouse with 35 mm lens
70 mm accessories bag
camcorder L1 bag
MIPS Laptop
Personal effects and hard-drives

In Priroda or available:

STS-84 resupply bag #18, 19, and 20 camera bags with 35 and 70 mm film
EVA Camera 35 mm with 35 mm lens
2 cassette tapes for camcorder
QUELD with 10 disks
MGBX in full; MiSDE in full
EDLS in full (11 WORM disks)
MIM
CGEL and has 2 video tapes being used at the moment
SAMS present with 1 disk unused and 2 in it
MIPS optical drive, controller, 2 laptops
Sleep Headband kits
Spare MIPS controller
BTS; GASMAP accessories
DCAM trays
Beetle
CPA and all items that came with it
Ethylene Glycol Real Time Sampling Kits
SiGB accessories
Greenhouse, everything but leaf bags
1 floppy for GH

Need:

10 floppies; 3 toothpaste; 3 toothbrushes
20 soaps (betadine); 4 shampoo bottles
1 crew notebook
Benzoyl peroxide; Tylenol - 100 tabs
Spare light bulbs approx. 20 (down to 3)
100 AA batteries
expanders
Therabands
can get by with EMK and MBK
1 pair of sunglasses
PCMCIA card coming from Houston (files from Home)

TEF was turned off ......loss of comm"
(the need list is being worked through the POSA/MOST, this is not action item to the community, just FYI)

  • submit to reddit


Loading







Join our mailing list




Monthly Archives

About this Entry

This page contains a single entry by Keith Cowing published on June 27, 1997 9:03 PM.

Hearing: Defining NASA's Mission and America's Vision for the Future of Space Exploration was the previous entry in this blog.

Charlie Harlin's Thoughts on Spaceflight Safety is the next entry in this blog.

Find recent content on the main index or look in the archives to find all content.